A bank's loan officer rates applicants for credit. The ratings are normally distributed with a mean of 200 and a standard deviation of 50. If an applicant is randomly selected, find the probability of a rating that is between 200 and 275. Write your answer as a decimal rounded to 4 places.
Answer #1

Solution :

Given that ,

mean = = 200

standard deviation = = 50

P(200 < x < 275) = P((200- 200)/ 50) < (x - ) /  < (275 -200) / 50) )

= P(0 < z < 1.5)

= P(z < 1.5) - P(z < 0)

= 0.9332 - 0.5000

= 0.4332

Probability = 0.4332

The probability of a rating that is between 200 and 275 is 0.4332